# Harmonic analysis of atomic vapor magnetization for in situ calibration of tri-axial time-dependent magnetic field

### v1

@@ Abstract
We introduce a methodology to calibrate in situ a set of coils generating bi- or tri-axial magnetic fields, at frequencies where a calibration performed in static condition would be inaccurate. The coil constants are determined in a two-step procedure. Considering the presence of a static and of a time-dependent field, firstly, the static one is oriented perpendicularly to the polarization plane of a time dependent one; secondly, the polarization of the latter is made accurately circular. The methodology uses harmonic analysis of one component of the magnetization of an atomic sample whose spins adiabatically follow the time-dependent field.
@@ I Introduction
The application of precisely assigned magnetic fields, with tailored spatial and temporal distribution is at the core of many precision experiments in modern physics and particularly in atomic laser spectroscopy and quantum optics [1, 2]. The generation of well controlled triaxial, time-dependent magnetic fields is of interest also for other application areas [3]. Solenoids, Helmholtz pairs and other specifically designed coil arrangements [4] are commonly used to achieve the desired field structure. Numerically controlled power supplies permit fine adjustments of the magnetic field components as well as the generation of custom-designed time-dependent magnetic fields.
@@
In instances requiring a two- or three-dimensional time-dependent field, setting its Cartesian and Fourier components with precise amplitude and relative phase may have a paramount importance [5, 6, 1].
The construction and characterization of magnetic field generators may follow both a priori and a posteriori approaches. The former is based on designing the current distribution (coil shape, and current waveform) in view of producing the appropriate fields, on the basis of the Maxwell’s laws. The a posteriori approach consists in measuring the actually produced field and to adjust its features with opportune additional coils or with opportune variations of the driving current(s). This second approach is often essential to improve the accuracy. In fact, the magnetic field generators are often surrounded by conducting materials or magnetic shields whose effects cannot be precisely taken into account in simulations or calculations.
@@
Several kinds of sensors can be used to measure the actual field, each of them coming with its inherent degree of precision, accuracy, spatial and time resolution, robustness etc. The use of Hall-effect or fluxgate sensors is often a favorite choice for their practicality, simple structure and vector response [7, 8]. However, in some applications they are not sufficiently precise or are not compatible with existing constraints.
@@
Laser spectroscopy with its unrivaled precision is an area of research that can require extreme accuracy in magnetic field control. At the same time, it enables the construction of very sensitive magnetic detectors, suitable for nicely performing such task.
@@
As proposed by Breschi et al.[9], an elegant and effective way to control the field with the accuracy required in atomic spectroscopy experiments and with detectors easily fitting in the setup is based on using the atomic sample itself as magnetometric sensor. This approach takes advantage from the high performance of optically pumped magnetometers, and avoids constraint issues because the atomic sample is already available as a part (typically the core!) of the experimental setup. This idea is at the basis of the in-situ calibration procedures that have been reported by several researchers.
@@
H.Zhang et al. [10] present a method to infer the coil constant (ratio between field and current) with a high precision by means of a high-performance hyperpolarized He magnetometer.
@@
A similar calibration method is proposed by Yao et al. [11], who use a hybrid potassium-rubidium magnetometer, analyzing the magnetic resonance under varying (swept) magnetic fields.
@@
Zhao et al. [13] use calibrated coils to obtain vector response from a self-oscillating rubidium optically-pumped magnetometer interrogated by far detuned (hence non-destructive) probe radiation. This approach has similarities with that proposed by G.Zhang et al. [14] using a forced (Bell and Bloom) magnetometer. In this case atoms are optically pumped by elliptically polarized radiation, which permits to achieve a vector response from the analysis of orientation and alignment dynamics.
@@
More recently Wang et al. [15] proposed a coil calibration method applied to a spin-exchange-relaxation-free optically pumped magnetometer based on the dynamic (transient) response of such system upon application of sudden variation of magnetic field along the three orthogonal directions.
@@
A calibration method based on magnetic induction detection with an accurately aligned pickup coil has been proposed by K.Zhang et al. [16] for Helmholtz coils operated in a frequency range from few tens to several hundreds Hz.
@@
A general feature shared by all the above mentioned works is to reduce the field measurement to a current (more commonly a voltage) measurement through an opportune set of calibration factors to be calculated or measured. The current-field linear dependence makes indeed possible to indirectly monitor the field by measuring the voltage across an assigned resistor connected in series with the field-generating coil. Those calibration factors are typically determined under static or quasi-static conditions, but can be used also in the case of time-dependent field, under the hypothesis that the instantaneous field can be inferred from the corresponding instantaneous voltage drop on the series resistor.
@@
The last assumption may be no longer valid in case of fast varying fields and currents. Indeed spurious phenomena (e.g. parasitic capacitance of cables and coils) can cause the measured current across the series resistor be different from that actually flowing in the coil, moreover unperceived eddy currents may be induced in the coil proximity. Thus the high frequency terms of the produced field may substantially differ, in amplitude and phase, from the corresponding terms inferred from the voltage drop measured on the monitor resistance [7].
In the recent years our research group has devoted a substantial effort to study the dynamics of atomic spins evolving in (or driven by) arbitrarily oriented time-dependent magnetic fields, with tailored sets of Fourier components applied along several spatial directions. In this research we have studied [17, 18] and applied [19] interesting features of phenomena commonly referred to as magnetic dressing [20], focusing on cases where the dressing field contains various Fourier and Cartesian components [21, 6]. These studies require an accurate control of the amplitude and phases of the magnetic field components applied to the atomic sample.
The core of the presented work is an innovative methodology aimed to achieve an accurate coil calibration over a broad frequency range, beyond the limit at which coil constants determined under static conditions become imprecise.
The proposed methodology makes use of a harmonic analysis of the signal that monitors one component of the atom polarization that evolves in the presence of a strong rotating field.
@@
We consider a magnetic field that is ideally made by a time-dependent component that rotates on a plane and a static one oriented perpendicularly to that plane and our goal is to detect imperfections of such field structure, which may consist in non-perfect perpendicularity of the static field, and in elliptical polarization of the time-dependent one.
The developed procedure is correspondingly made of two steps. The first step is to orient the static field perpendicularly to the polarization plane of the rotating one. The second one serves to establish a circular polarization of the rotating field. This permits to identify and counteract possible spurious effects occurring at higher frequencies and affecting amplitudes and phases of different field components.
@@
The atomic sample is optically interrogated, with polarimetric techniques previously developed for magnetometric measurements, in such a way to produce a signal proportional to one component of the macroscopic magnetization. Diverse harmonic components of that signal bring information about the mentioned field imperfections, thus an inherently in-situ calibration procedure is developed on the basis of Fourier analysis.
@@ II Setup
The setup built to prepare and interrogate the atomic magnetization is described in Ref.[22], where it was used for magnetometric measurements. Briefly, Cs atoms are pumped by means of circularly polarized radiation tuned to the D1 line and probed with a weak linearly polarized radiation detuned by few GHz from the D2 line. A balanced polarimeter is used to measure the Faraday rotation of the probe polarization plane, which provides a signal proportional to the sample magnetization component along the probe-beam axis: this is the signal to be analyzed for our methodology. The two beams co-propagate along the $x$ direction and in magnetometric application a static field is applied along the $z$ direction. The balanced polarimeter uses a transimpedance amplifier with a bandwidth of several kHz, enabling the detection of signals with a spectrum in the audio range.
@@
The experiment is run in an unshielded environment, where the magnetic field and its first-order gradients are compensated by three large-size (180 cm) Helmholtz pairs and five quadrupole sources, all of them driven by numerically controlled current generators that can be set manually or automatically [23]. When needed, one of the field components is compensated partially to establish a static field.
@@
Additional coils enable the application of time-dependent fields along the three directions. The co-propagating arrangement of the pump- and probe-beams permits the use of a solenoid for the $x$ component, while the $y$ and $z$ components are generated by small size (45 cm and 50 cm size, respectively) squared Helmholtz pairs, each with 50 turns per coil. These three alternating field coils are supplied via audio-amplifiers that amplify three arbitrary waveforms generated by a digital-to-analog (DAC) converters card (NI 6343), as sketched in Fig.1. Depending on the required frequencies, matching series-impedances can be applied to improve the coupling between each amplifier and the respective coil. The current actually flowing in each coil is monitored by recording the voltage drop over series resistors, using analog-to-digital converters (ADC) available in the same NI-6343 card.
@@
Figure 1: Schematics of the time-dependent field generator and monitor for one Cartesian component of the time dependent field: DAC: digital-to-analog converter, programmed for arbitrary waveform generation; A: amplifier; MI: matching impedance; C: coil (solenoid or Helmholtz pair); MR: monitor resistance; ADC: analog to digital converter.
From the size, shape and number of turns in each coil it is possible to evaluate a priori the current-field calibration factor on the basis of basic equations, such as Biot-Savart and Ampère laws.
@@
An a posteriori calibration is then performed, under static conditions, with the help of the magnetometer, by measuring the variations of the Larmor frequency in response to changes of the current flowing in the coils. The Larmor frequency can be evaluated scanning the frequency of the pump modulation across the magnetic resonance as described in [22] or estimating the frequency of a free precession signal as described, e.g., in [24, 25].
@@
In the case of free-precession measurement, the procedure implemented for our setup is as follows: the pump laser is tuned to the $F_{g}=3\rightarrow F_{e}=4$ D1 transition for 100 ms, then is abruptly blue-detuned by about 30 GHz by reducing its drive current. As soon as the pump radiation is made off-resonant, the data acquisition starts, lasting other 100 ms as to measure the transient signal produced by the free precession of the atomic spins.
Refer to captionRefer to caption
Figure 2: Free precession signal in the time (a) and the frequency (b) domains.
@@
An example of such transient signal is reported in Fig.2. No appreciable harmonic distortion is caused at the photo-detection stage: no harmonic peaks emerge from the noise floor, indicating that they are more than 45 dB weaker than the fundamental tone. The localization of the fundamental tone peak in Fig.2b permits to evaluate the static field. In this case the magnetic field is estimated to be $1.071\mu$T ±$\pm 3$nT. A consistent value ($1.070\mu$T ±$\pm 3$nT) was obtained from the analysis of the resonance profile obtained in forced conditions, i.e. running the magnetometer in the Bell and Bloom configuration [22]. The ±$\pm 3$nT uncertainties are estimated from the standard deviation over large measurement sets, and they are due to the ambient field fluctuations.
@@
As a starting point, the generation of the time-dependent field from each coil uses the calibration factors determined under static conditions. To this aim, the DAC is programmed to generate a signal with known amplitude and phase at the desired frequency; this signal is amplified and applied to the coil; the voltage drop on the monitor resistor is acquired and analyzed to extract its phase and amplitude; the latter is then converted to current and then to field on the basis of the monitor resistance and the static coil constant; finally the DAC output is scaled and rephased to achieve the field settings (phase and amplitude) set by the operator. Being DAC and ADC of the diverse coils synchronous, this procedure enables the generation of field components with relative phase and amplitudes assigned on the basis of the static calibration factors.
@@
Transient signal measurements are also applied to characterize time-dependent, three-dimensional magnetic fields, i.e. for the dynamic calibration procedure at the focus of the present work. In this case the signal is not due to a free precession of the spins around a static field, but to their strongly driven dynamics. The latter consists in an adiabatic following of the applied time-dependent field, as discussed in the following (Sec.III). The recorded signal consists of a damped oscillation as in the case of free induction decay measurement, but here the oscillation frequency is that of the rotating field.
@@ III Methodology
Our aim is an accurate determination of the coil constants for two orthogonal field generators, taking into account the amplitude and phase deviations that occur when the frequency of the field to be generated is too high, such as to make the constants determined in static conditions imprecise. To this aim, we generate a rotating field with a nominally circular polarization, using the static constants; we evaluate its actual ellipticity; and we eventually determine the corrections needed to compensate the detected discrepancies.
The idea at the basis of the proposed methodology relies on the adiabatic precession of the atomic sample magnetization around the instantaneous magnetic field to be characterized. The atomic spins follow adiabatically the time-dependent field under the condition that the instantaneous precession is much faster than the reorientation of the field. In other terms, with a time dependent field that rotates at an angular speed $\omega$, it is required that
@@
where $\gamma$ is the gyromagnetic factor and $B$ is the modulus of the field. We consider a field nominally made of a static component along $z$ and a rotating one on the $xy$ plane. Possible imperfections are considered, which may concern the orientation of the static field and some degree of ellipticity in the polarization of the time-dependent terms. Namely, we consider
@@
with $m_{1},m_{2}$ accounting for terms that make the static field not perpendicular to the polarization plane $xy$, and $\varepsilon,\phi$ to describe imperfections of the time-dependent field polarization. $B_{0}$ is the nominal amplitude of the rotating field and $m_{3}B_{0}$ is the nominal amplitude of the static one.
The aim is to obtain a time dependent field with a circular polarization on a plane perpendicular to the static term, i.e. the achiemement of $m_{1}=m_{2}=\varepsilon=\phi=0$ condition (see eq.(2)).
As known, in the hypothesis of adiabatic following, the macroscopic magnetization $\vec{M}$ maintains a constant modulus $M$ (apart from the decay due to the relaxation mechanisms) while precessing around $\vec{B}(t)$.
Among the components of $\vec{M}$, the one parallel to $\vec{B}$ remains approximately constant, while the perpendicular ones oscillate at high frequency, whose instantaneous value is $\gamma|B|$.
@@
The developed procedure is made of a first step aimed to make $m_{1}=m_{2}=0$, and a second one to make $\varepsilon=\phi=0$. Both the scopes can be pursued on the basis of Fourier analysis of the Taylor approximation of the quantity $M_{\parallel x}(t)$ (eq.(3)), for which the relevant details are reported in the Appendix.
III.1 Alignment of the static field along $z$
A straightforward procedure to orient the static field along $z$ is based on minimizing of the Larmor frequency by varying the currents that drive the $x$ and $y$ coils (in the absence of time-dependent fields): the minimum is achieved when the $x$ and $y$ components of the field are fully compensated.
In the presence of a rotating field on the $xy$ plane, a harmonic analysis of the detected signal (eq.(A.1)) can be used to reveal the presence of spurious static components on the $xy$ plane, as well.

### v2

@@ Abstract
We introduce a methodology to calibrate in situ a set of coils generating bi- or tri-axial magnetic fields, at frequencies where a calibration performed under static conditions would be inaccurate. The coil constants are determined in a two-step procedure. Considering the presence of a static and a time-dependent field, firstly, the static one is oriented perpendicularly to the polarization plane of a time-dependent one; secondly, the polarization of the latter is made accurately circular. The methodology uses harmonic analysis of one component of the magnetization of an atomic sample whose spins adiabatically follow the time-dependent field.
@@ I Introduction
The application of precisely assigned magnetic fields with tailored spatial and temporal distribution is at the core of many precision experiments in modern physics particularly in atomic laser spectroscopy and quantum optics [1, 2]. Generating well-controlled triaxial, time-dependent magnetic fields is of interest also for other application areas [3]. Solenoids, Helmholtz pairs, and other specifically designed coil arrangements [4] are commonly used to achieve the desired field structure. Numerically controlled power supplies enable precise adjustments of the magnetic field components and the generation of custom-designed time-dependent magnetic fields.
@@
In instances requiring a two- or three-dimensional time-dependent field, setting its Cartesian and Fourier components with accurate amplitude and relative phase is of paramount importance [5, 6, 1].
The works cited above consider time-dependent fields in the 100 Hz - 20 kHz range. It is worth noting that other research requires bi- or tri-axial fields at much higher frequencies, up to the radio [7] and microwave [8, 9, 10] ranges. We consider a methodology based on atomic spins that adiabatically follow the time-dependent field. This requirement limits the frequency at which the methodology can be practically applied to the audio range.
@@
The construction and characterization of magnetic field generators can follow both a priori and a posteriori approaches. The former involves designing the current distribution (coil shape and current waveform) aiming to produce the appropriate fields based on Maxwell’s laws. The a posteriori approach involves measuring the actual field and adjusting it using additional coils or varying the driving current(s). This can be essential for accuracy because the magnetic field generators are often surrounded by conducting materials or magnetic shields whose effects cannot be precisely considered in simulations or calculations. Several kinds of sensors can be used to measure the actual field, each of them coming with its inherent degree of precision, accuracy, spatial and time resolution, robustness etc. The use of Hall-effect or fluxgate sensors is often a favorite choice for their practicality, simple structure, and vector response [11, 12]. However, in some applications, they are not sufficiently precise or are not compatible with existing constraints.
@@
Laser spectroscopy with its unrivaled precision is an area of research that can require extreme accuracy in magnetic field control. At the same time, it enables the construction of excellent magnetic detectors.
@@
As proposed by Breschi et al.[13], an elegant and effective way to control the field with the required accuracy and with detectors easily fitting in atomic spectroscopy setups is based on using the atomic sample itself as a magnetometric sensor. This concept forms the basis of in-situ calibration procedures reported by several researchers.
@@
H.Zhang et al. [14] present a method to infer the coil constant (ratio between field and current) with high precision employing a high-performance hyperpolarized He magnetometer.
@@
A similar calibration method is proposed by Yao et al. [15], who use a hybrid potassium-rubidium magnetometer and analyze the magnetic resonance under varying (swept) magnetic fields.
@@
Zhao et al. [17] use calibrated coils to obtain vector response from a self-oscillating rubidium optically-pumped magnetometer non-destructively interrogated by far detuned probe radiation. This approach has similarities with that proposed by G.Zhang et al. [18], which uses a forced (Bell and Bloom) magnetometer. In this case, atoms are optically pumped by elliptically polarized radiation, which permits the achievement of a vector response from the analysis of orientation and alignment dynamics.
@@
More recently K.Wang et al. [19] proposed a coil calibration method applied to a spin-exchange-relaxation-free optically-pumped magnetometer based on the dynamic (transient) response of such system upon application of sudden variation of the magnetic field along the three orthogonal directions.
@@
A calibration method based on magnetic induction detection with an accurately aligned pickup coil has been proposed by K.Zhang et al. [20] for Helmholtz coils operated in a frequency range from few tens to several hundred Hz.
@@
A general feature shared by all the above-mentioned works is to reduce the field measurement to a current (more commonly a voltage) measurement through an opportune set of calibration factors to be calculated or measured. The current-field linear dependence makes it possible to indirectly monitor the field by measuring the voltage across an assigned resistor connected in series with the field-generating coil. Those calibration factors are typically determined under static or quasi-static conditions but can be used also in the case of a time-dependent field, under the hypothesis that the instantaneous field can be inferred from the corresponding instantaneous voltage drop on the series resistor.
@@
The last assumption may be no longer valid in case of fast varying fields and currents. Spurious phenomena (e.g. parasitic capacitance of cables and coils) can cause the measured current across the series resistor to be different from that actually flowing in the coil, moreover, unperceived eddy currents may be induced in the coil proximity. Thus the high-frequency terms of the produced field may substantially differ, in amplitude and phase, from those inferred from the voltage drop measured on the monitor resistance [11, 1].
The motivation that inspired the present work is the interest in studying the dynamics of atomic spins that evolve in (or are driven by) arbitrarily oriented time-dependent magnetic fields. In this research we have studied [21, 22] and applied [23] phenomena commonly referred to as magnetic dressing [24], focusing on cases where the dressing field contains various Fourier and Cartesian components [25, 6]. This research requires a very accurate control of the amplitude and relative phases of the magnetic field components. We present a method to accurately generate a circularly polarized field. This method permits an accurate a posteriori determination of the calibration parameters, which can subsequently be used to produce any kind of field polarization, over a broad frequency range, beyond the limit at which coil constants determined under static conditions become imprecise. The achieved results will find application in all cases where oscillating field components along diverse axes must be generated with very accurate relative phases and amplitudes.
@@
We consider a magnetic field ideally made of a time-dependent component that rotates on a plane and a static one oriented perpendicularly to that plane. Our goal is to detect imperfections of such field structure, which may consist of non-perfect perpendicularity of the static field and of elliptical polarization of the time-dependent one.
The developed procedure is correspondingly made of two steps. The first makes the static field perpendicular to the polarization plane of the time-dependent one. The second one establishes a circular polarization of the latter.
@@
An optically-pumped atomic sample is interrogated employing polarimetric techniques, to produce a signal proportional to one component of the macroscopic magnetization. Diverse harmonic components of this signal bring information about the mentioned field imperfections, allowing an inherently in-situ calibration procedure to be developed based on Fourier analysis.
@@ II Setup
The setup built to prepare and interrogate the atomic magnetization is described in Ref.[26]. Briefly (see Fig.1), Cs atoms are pumped employing circularly polarized radiation (at mW level) tuned to the D1 line and probed with a weak ($\mu$W level) linearly polarized radiation detuned by a few GHz from the D2 line. The D1 radiation is brought in resonance for tens of milliseconds and suddenly detuned to start the measurement. A balanced polarimeter is used to detect the Faraday rotation of the probe polarization plane, providing a signal proportional to the sample magnetization component along the probe-beam axis. The two beams co-propagate along the $x$ direction and in magnetometric application a static field is applied along the $z$ direction. The balanced polarimeter uses a transimpedance amplifier with a bandwidth of several kHz, enabling the detection of signals within the audio spectral range.
Figure 1: Schematics of the setup. Cs atoms contained in a sealed cell (Cs) are optically pumped by circularly polarized radiation (D1) and probed by linearly polarized radiation (D2). The two laser beams co-propagate along $x$ and an interferential filter (IF) blocks the pump before the polarimeter (POL) that detects the Faraday rotation of the probe polarization. Three large-size Helmholtz pairs control the static magnetic field (only two of them, $B_{DCz}$ and $B_{DCx}$, are shown). Two smaller Helmholtz pairs ($B_{DCy}$ and $B_{DCz}$) generate time-dependent components along $y$ and $z$, while a solenoid ($B_{ACx}$) is used for the $x$. Several quadrupoles (not shown) help counteract static inhomogeneities.
@@
The experiment is run in an unshielded environment, where the magnetic field and its first-order gradients are controlled by three large-size (180 cm) Helmholtz pairs and five quadrupole sources, all of them driven by numerically controlled direct current (DC) generators that can be set manually or automatically [27].
@@
Additional coils enable the application of time-dependent (AC) fields along the three directions, two of which are used in this experiment. The co-propagating arrangement of the pump- and probe-beams permits the use of a solenoid for the $x$ component, while the $y$ and $z$ components are generated by small size (45 cm and 50 cm size, respectively) squared Helmholtz pairs, each with 50 turns per coil. These AC coils are supplied via audio-amplifiers that amplify arbitrary waveforms generated by a digital-to-analog (DAC) converters card (NI 6343), as sketched in Fig.2. Depending on the required frequencies, matching series-impedances can be applied to improve the coupling between each amplifier and the respective coil. The current flowing in each coil is monitored by recording the voltage drop over series resistors, using analog-to-digital converters (ADC) available in the same NI-6343 card.
@@
Figure 2: Schematics of the time-dependent field generator and monitor for one Cartesian component of the time-dependent field: DAC: digital-to-analog converter, programmed for arbitrary waveform generation; A: amplifier; MI: matching impedance; C: coil (solenoid or Helmholtz pair); MR: monitor resistance; ADC: analog to digital converter.
@@
An a priori calibration of all the coils is performed based on size, shape and number of turns using the Biot-Savart law. An a posteriori calibration is then performed, under static conditions, by measuring the Larmor frequency in response to assigned DC currents applied to the coils. The Larmor frequency can be evaluated by scanning the frequency of the pump modulation across the magnetic resonance as described in [26] or estimating the frequency of a free precession signal (FPS) as described, e.g., in [28, 29].
@@
In the case of FPS measurements, the procedure implemented is as follows: the pump diode-laser is tuned to the $F_{g}=3\rightarrow F_{e}=4$ D1 transition for 100 ms, then is abruptly blue-detuned by about 30 GHz by reducing its drive current. As soon as the pump radiation is made off-resonant, the data acquisition starts, lasting another 100 ms.
Figure 3: FPS in the time (a) and the frequency (b) domains. The red, blue, and orange lines indicate the precession frequency and its $2^{\mathrm{nd}}$, an $3^{\mathrm{rd}}$ harmonics, respectively.
@@
An example of such FPS is reported in Fig.3. No appreciable harmonic distortion is observed: no harmonic peaks emerge from the noise floor, confirming the linearity of the photo-detection stage. The localization of the fundamental tone peak in Fig.3b permits the evaluation of the static field. In this case, the magnetic field is estimated to be $1.071\mu$T ±$\pm 3$nT. A consistent value ($1.070\mu$T ±$\pm 3$nT) was obtained from the analysis of the resonance profile obtained in forced conditions, i.e. running the magnetometer in the Bell and Bloom configuration [26]. The ±$\pm 3$nT uncertainties are estimated from the standard deviation over large measurement sets, and they are due to the ambient field fluctuations.
@@
The AC field generated by each coil is designed based on the calibration factors determined under static conditions. To this aim, the DAC is programmed to generate a signal with known amplitude and phase at the desired frequency; this signal is amplified and applied to the coil; the voltage drop on the monitor resistor is acquired and analyzed to extract its phase and amplitude; the latter is then converted to current and then to field based on the monitor resistance and the static coil constant; finally the DAC output is scaled and rephased to achieve the field settings (phase and amplitude) set by the operator. Being DAC and ADC of the diverse coils synchronous, this procedure enables the generation of field components with assigned relative phase and amplitudes but relies on the static calibration factors.
@@
Transient signal measurements are also applied to characterize time-dependent, two-dimensional magnetic fields, i.e. for the dynamic calibration procedure at the focus of the present work. In this case, the signal is not a FPS but originates from atomic spins that adiabatically follow the time-dependent applied field. In both cases, the recorded signal is a damped oscillation but in this second case the oscillation frequency is that of the rotating field.
@@ III Methodology
The proposed methodology requires that the atomic magnetization adiabatically follows the magnetic field to be characterized, i.e., the atomic spins must precess much faster than changes in field orientation. With a time-dependent field that rotates at an angular speed $\omega$, it is required that
@@
where $\gamma$ is the gyromagnetic factor and $B$ is the modulus of the field.
We consider a field nominally made of a static component along $z$ and a rotating one on the $xy$ plane. Possible imperfections are considered, which may concern the orientation of the DC field and some degree of ellipticity in the polarization of the time-dependent field. Namely, we consider
@@
with $m_{1},m_{2}$ accounting for spurious static field components on the polarization plane $xy$, and $\varepsilon,\phi$ for imperfections of the time-dependent field polarization. $B_{0}$ is the nominal amplitude of the rotating field and $m_{3}B_{0}$ is the nominal amplitude of the static one.
The aim is to obtain a time-dependent field with a circular polarization, i.e $\varepsilon=\phi=0$, on a plane perpendicular to the static term, i.e. $m_{1}=m_{2}=0$.
In the hypothesis of adiabatic following, the macroscopic magnetization $\vec{M}$ maintains a constant modulus $M$ (apart from the decay due to the relaxation mechanisms) while precessing around $\vec{B}(t)$. Among the components of $\vec{M}$, the one parallel to $\vec{B}$ remains approximately constant, while the perpendicular ones oscillate at the high frequency $\gamma|B|$ [30].
@@
The developed procedure is made of a first step aimed to make $m_{1}=m_{2}=0$, and a second one to make $\varepsilon=\phi=0$. Both the scopes can be pursued on the basis of Fourier analysis of the Taylor approximation of the quantity $M_{\parallel x}(t)$ (eq.(3)), the Appendix reports the relevant details.
III.1 Alignment of the static field along $z$ ($m_{1}=m_{2}=0$ condition)
Traditionally, the procedure to orient a static field along an axis (let it be $z$, as in our case) is based on minimizing the Larmor frequency by varying the currents that drive the $x$ and $y$ coils: the minimum is achieved when the $x$ and $y$ components of the field are fully compensated.
In the presence of a rotating field on the $xy$ plane, harmonic analysis of the detected signal (eq.(A.1)) can be used to reveal the presence of static components on the $xy$ plane, as well.